OpenStudy (anonymous):

Find the solutions for each equation on [0,2pi). 1. 2sin^2Ө=1-sinӨ 2. 2tanӨ-sec^2Ө=0 3. sin2Ө+sinӨ=0

OpenStudy (psymon):

Well let's start with the first one. This is a quadratic equation, but instead of x we have sine. So we want to factor this just liek a quadratic, which means we set everything equal to 0 and factor: \[2\sin ^{2}\theta+\sin \theta-1=0\]Now if you were to pretend that sin(x) was just x, would you know how to factor that quadratic?

OpenStudy (anonymous):

we are pretending the sinӨ is x?

OpenStudy (psymon):

Correct. Because this is done EXACTLY like if we had \[2x ^{2}+x -1 = 0\]we factor it the same way, its just its sinx instead of x.

OpenStudy (anonymous):

(2x-1)(x+1)=0 got it whats next?

OpenStudy (psymon):

Now we just need to replace x with sin x and set both factors equal to 0. So we have this: \[2\sin \theta-1 = 0\]and \[\sin \theta +1 = 0\] Now we first solve for sin(theta) then solve for the actual theta.

OpenStudy (anonymous):

and you do that for both? So sinӨ=1/2 and sinӨ= -1

OpenStudy (psymon):

Correct. Now do you know the values of the unit circle or have a unit circle chart handy?

OpenStudy (anonymous):

yes how do I used that?

OpenStudy (psymon):

Right. Well, the unit circle chart will tell you where sin = 1/2 and where sin = -1. If you have a way to look it up then you can find which angles by looking at the chart.

OpenStudy (anonymous):

and the angles are my answer?

OpenStudy (psymon):

Correct. You will have two angles that equal 1/2 and one angle that = -1. So you will have 3 answers.

OpenStudy (anonymous):

okay thanks may need help with other two but don't know quite yet. :)

OpenStudy (psymon):

Did you find the correct angles?

OpenStudy (anonymous):

When looking at the circle is the 1/2 and -1 in the "x" or "y" part of the coordinate

OpenStudy (psymon):

Sin refers to the y-coordinate.

OpenStudy (anonymous):

30, 270, and 350 then

OpenStudy (psymon):

150 instead of 350.

OpenStudy (anonymous):

okay

OpenStudy (psymon):

Now the second one requires us to use an identity. Do you know of this identity? \[\tan ^{2}\theta+1 = \sec ^{2}\theta \]?

OpenStudy (anonymous):

uhm yes I don't remember how to use it tho?

OpenStudy (psymon):

Well our second problem is also a quadratic. The problem is we have a tangent and a secant. In order to do this problem, they must both be the same trig function. So we use this identity to change secant into tangent. Now we have: \[2\tan \theta -\sec ^{2}\theta=0\]But look what the identity above says. It says secant squared is the same as tangent squared + 1. So we have to substitute that.

OpenStudy (anonymous):

okay

OpenStudy (psymon):

Now just make sure to subsitute it properly, not making sign mistakes: \[2\tan \theta -(\tan ^{2}\theta +1)=0\]Now would you know how to solve that in the same way we did with the first problem?

OpenStudy (anonymous):

yes I would thanks

OpenStudy (psymon):

Alright, cool. So we would end up with \[\tan \theta = 1\]Now to solve this one, we need to know that: \[\tan \theta = \frac{ \sin \theta }{ \cos \theta }\]So when we check our unit circle, we want to know when: \[\frac{ \sin \theta }{ \cos \theta }=1 \]There will be two answers :3

OpenStudy (psymon):

If you got the answer to this one we can go to the last one.

OpenStudy (anonymous):

no sorry I didn't get that one yet

OpenStudy (psymon):

Ah, okay. Well, we need to find angles where if we divide sin and cos we get 1. This means that the values for sin and cosine must be the same at the angles we need. There are two spots where this is true.

OpenStudy (anonymous):

90 and 360

OpenStudy (psymon):

Well, at 90 degrees cosine is 0 and at 360 sin is 0, so theres no way that they can divide and get you 1. In fact, tangent is completely undefined at 90.

OpenStudy (anonymous):

so am I looking for sin and cosine at 1 or tangent?

OpenStudy (psymon):

You want a tangent that is 1. But to get a tangent that is 1, you need to find out where sin and cosine are the exact same value. Since tangent = sin/cos, the only way two numbers can divide and equal 1 is if they are the exact same number.

OpenStudy (anonymous):

do you have a unit circle handy tat you could help me? I'm looking at this one and just cant figure it out

OpenStudy (psymon):

Im sure we can google and find one. I just have the values memorized personally. Um...lemme look.

OpenStudy (anonymous):

okay thanks:)

OpenStudy (psymon):

There are two spots where sine and cosine are the exact same. And remember that by exact same, that means the same sign as well.

OpenStudy (anonymous):

wait im writing out my work and can we go back to where we switched sec^2 theta to tan^2+1 what happens after I do that step im sorry

OpenStudy (psymon):

Ah, okay. Well, we had: \[2\tan \theta - (\tan ^{2}\theta + 1) = 0\]Set everything equal to 0 and make sure the tan^2 term is positive: So we'll add the quantity of tan^2 + 1 to the other side of the equation and then move 2tan to the other side, too: So we end up with: \[\tan ^{2}\theta - 2\tan \theta +1 = 0\] Now we just factor this liek a quadratic, same way we did in the first problem.

OpenStudy (anonymous):

and could u write out the equations so I can see if I'm correct?

OpenStudy (psymon):

Factors of 1 that add up to negative 2 = (-1)(-1). Therefore we factor this into: \[(\tan \theta -1)^{2}=0\]The two factors are the same,so we set one of them equal to 0: \[\tan \theta =1 \]Now we find what angles havea tangent value of 1. This is done by checking our unit circle and seeing where sine and cosine are the exact same values. THis occurs at two angles.

OpenStudy (anonymous):

and both sine and cosine are the y coordinate?

OpenStudy (psymon):

Cosine is the x-coordinate and sine is the y-coordinate.

OpenStudy (anonymous):

is it 60 and 30?

OpenStudy (psymon):

Neither. at 30 degrees, sin is 1/2 and cosine is sqrt(3)/2. At 60 degrees it is flip-flopped. sine is sqrt(3)/2 and cosine is 1/2. We want sine and cosine to be the exact same values. That link I posted of the unti circle should show you.

OpenStudy (anonymous):

45 and 225?

OpenStudy (psymon):

There you go xD

OpenStudy (anonymous):

yay!! can we do the last one? :)

OpenStudy (psymon):

Yeah, sure. Now this one will be a little bit different. We do have everything sine, but we have different angles. One is 2theta, the other is just theta. We cannot have that, these must be the same angle. So this means we have to use the double-angle formula on the sin2(theta). Would you happen to know it?

OpenStudy (anonymous):

nope LOL but if you do just show me the work I can usually teach myself

OpenStudy (psymon):

Well, the double-angle formula for sin2(theta) is: \[2\sin \theta \cos \theta \] So this means we rewrite what we have as: \[2\sin \theta \cos \theta + \sin \theta = 0\]So now we just have to factor this the best we can, which isn't much, but you see how to do it?

OpenStudy (anonymous):

yes but now how would you factor that mess LOL?

OpenStudy (psymon):

Only one thing you can do and thats factor out a sin xD

OpenStudy (anonymous):

so sin(1/cosӨ)=0 and sinӨ=0?

OpenStudy (psymon):

Nope: Like this: \[\sin \theta(2\cos \theta +1)=0 \]can you see how?

OpenStudy (anonymous):

so im only gonna have one function?

OpenStudy (psymon):

Nope, you have 2. sin(theta) counts as a factor, too. You set both factors to 0 like this: \[\sin \theta = 0\] \[2\cos \theta + 1 = 0\]

OpenStudy (anonymous):

then so Ill end up with 0 and -1/2 right?

OpenStudy (anonymous):

0 being sine and -1/2 being cosine

OpenStudy (psymon):

Thats what theyll be equal to. Now you need to solve for theta in both equations.

OpenStudy (anonymous):

by the unit circle right? how many angles should I get

OpenStudy (psymon):

4 : ) 2 for each.

OpenStudy (anonymous):

240, 120, 180, and 360?

OpenStudy (psymon):

Sounds good xDD

OpenStudy (anonymous):

YES!!! Thank soooo much:) I really appreciate it

OpenStudy (psymon):

yep, no problem :3